OpenStudy (anonymous):

Help please! Indicate the equation of the given line in standard form. The line that is the perpendicular bisector of the segment whose endpoints are R(-1, 6) and S(5, 5)

hartnn (hartnn):

can u find midpoint of R and S ?

OpenStudy (anonymous):

I'm not sure how to.

hartnn (hartnn):

mid-point of (x1,y1) and (x2,y2) is x= (x1+x2)/2 y= (y1+y2)/2

OpenStudy (anonymous):

What do I do after that?

hartnn (hartnn):

what is the mid-point u get ?

OpenStudy (anonymous):

2.5, 5

hartnn (hartnn):

5-1=4 4/2=2 6+5=11 11/2=5.5 so the midpoint is (2,5.5) ok?

OpenStudy (anonymous):

Oh, gotcha. Sorry, I read that backwards in my head. Haha what do I do after that, to get the equation of the line?

hartnn (hartnn):

now find the slope of given line: The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

I got - 1/7

hartnn (hartnn):

how ?

OpenStudy (anonymous):

6 - 5 = 1 -1 - 5 = -6 I typed 7 by mistake instead of 6.

hartnn (hartnn):

so its -1/6 then slope of perpendicular will be reciprocal inverse. so it'll be m=6 now you have slope and point (2,5.5) can u find the equation of that line ?

OpenStudy (anonymous):

y-2=6(x-5.5) ?

hartnn (hartnn):

u reversed the points. y-y1=m(x-x1) y-5.5 = 6(x-2)

OpenStudy (anonymous):

Got it. Thanks!

hartnn (hartnn):

welcome ^_^